Find equation for
(-1/2,-3/2) and slope of -2/3
Pelase show steps

Answer by rothauserc(4718)

the slope intercept form of a line is y = mx +b where m is slope and b is the y intercept (where the line crosses the y axis)
:
m = -2/3, then
y = -2x/3 + b
:
we find b by substituting the x and y coordinates of our point
-3/2 = (-2/3)(-1/2) + b
-3/2 = 2/6 + b
note that -3/2 is -9/6
-9/6 = 2/6 + b
b = -11/6
:
therefore we have,
y = -2x/3 - 11/6
:
our equation of the line is
f(x) = -2x/3 - 11/6
: